### v3

# GaMeS: Mesh-Based Adapting and Modification of Gaussian Splatting

Abstract
Recently, a range of neural network-based methods for image rendering have been introduced. One such widely-researched neural radiance field (NeRF) relies on a neural network to represent 3D scenes, allowing for realistic view synthesis from a small number of 2D images. However, most NeRF models are constrained by long training and inference times. In comparison, Gaussian Splatting (GS) is a novel, state-of-the-art technique for rendering points in a 3D scene by approximating their contribution to image pixels through Gaussian distributions, warranting fast training and swift, real-time rendering. A drawback of GS is the absence of a well-defined approach for its conditioning due to the necessity to condition several hundred thousand Gaussian components. To solve this, we introduce the Gaussian Mesh Splatting (GaMeS) model, which allows modification of Gaussian components in a similar way as meshes. We parameterize each Gaussian component by the vertices of the mesh face. Furthermore, our model needs mesh initialization on input or estimated mesh during training. We also define Gaussian splats solely based on their location on the mesh, allowing for automatic adjustments in position, scale, and rotation during animation. As a result, we obtain a real-time rendering of editable GS.
Gaussian Splatting, Mesh, NeRF, Rendering
Before modification   After modification
Figure 1: GaMeS produce a hybrid of Gaussian Splating and mesh representations. Therefore, GaMeS allows real-time modification and adaptation of Gaussian Splatting. The supplementary material provides more examples, including a video illustrating our results.
1 Introduction
Recently, we have observed the emergence of several promising methods for rendering unseen views of 3D objects and scenes using neural networks. For instance, Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) have rapidly grown in popularity within the computer vision and graphics communities (Gao et al., 2022) as they enable the creation of high-quality renders. Despite this interest and growing body of related research, the long training and inference time remains an unsolved challenge for NeRFs.
In contrast, latterly introduced Gaussian Splatting (GS) (Kerbl et al., 2023) offers swift training and real-time rendering capabilities. What is unique to this method is that it represents 3D objects using Gaussian distributions (i.e. Gaussians). Hence, it does not rely on any neural network. Consequently, Gaussians can be employed in a manner akin to manipulating 3D point clouds or meshes, allowing for actions like resizing or repositioning in 3D space. Nonetheless, practical challenges may arise when altering Gaussian positions, particularly in accurately tracking changes in the shape of Gaussian components, such as the ellipses. Moreover, scaling Gaussian components proves challenging when the object undergoes resizing, which is not an issue for classical meshes, as their triangle faces can be readily updated when vertex positions are adjusted.
The above constraints may be resolved by constructing Gaussian directly on the mesh, as shown by SuGaR (Guédon & Lepetit, 2023). Here, the authors introduced a regularization term in the Gaussian splat cost function to promote optimal alignment of the Gaussians with the scene’s surface. The SuGaR model uses signed distance functions (SDF) from the vanilla GS and minimizes the difference between this SDF and its actual value computed for the Gaussians. Moreover, the authors propose a method for extracting mesh from the GS model and introduce an optional refinement strategy that binds Gaussians to the surface of the mesh and jointly optimizes these Gaussians and the mesh through GS rendering.
Figure 2: GaMeS can be effectively trained on large scenes to allow their modifications while preserving high-quality renders.
In our paper, we propose an alternative approach to such a task. Unlike SuGaR, we do not have two expensive stages dedicated to producing high-quality mesh. Instead, our Gaussian Mesh Splatting (GaMeS) uses existing mesh or approximates it during one-stage training. Our estimated mesh uses a set of triangle faces, which are not connected; therefore, we called it pseudo-mesh. Our pseudo-mesh is dedicated to modifying the GS model, not approximating objects’ surfaces. Implementing GaMeS involves positioning Gaussian components on the faces of the pseudo-mesh, ensuring proper alignment of these components with the mesh structure. Using our method, we can obtain comparable to state-of-the-art high-quality outcomes that can be attained for static scenes akin to the GS method. Moreover, any alterations to the mesh will automatically propagate updates to the corresponding Gaussian components, enabling real-time animation (see Fig. 1). Such an approach can also be applied in scenarios with preexisting mesh (see Fig. 2). Most importantly, we can train GS and then generate pseudo-mesh from Gaussian components to force a model to be editable. In such a strategy, our pseudo-mesh is not dedicated to describing the object’s surface but to editing the Gaussian Splatting model.
In summary, this work makes the following contributions:
  • We introduce a hybrid representation for 3D objects, seamlessly integrating mesh and GS.
  • GaMeS work on the simple pseudo-mesh that can be estimated directly from Gaussian components obtained in a single training without an expensive preprocessing strategy.
  • Our method relies only upon essential vector operation. Consequently, we can render dynamic scenes at a similar pace to their static counterpart.
2 Related Works
Point-based Gaussian representations have found a large number of potential application scenarios, including substituting point-cloud data (Eckart et al., 2016), molecular structures modeling (Blinn, 1982), or shape reconstruction (Keselman & Hebert, 2023). These representations can also be utilized in applications involving shadow (Nulkar & Mueller, 2001) and cloud rendering (Man, 2006). In addition, a new technique of 3D Gaussian Splatting (3D-GS) was recently introduced (Kerbl et al., 2023). This method couples splatting methods with point-based rendering for real-time rendering speed. The rendering quality of 3D-GS is comparable to that of Mip-NeRF (Barron et al., 2021), one of the best multilayer perceptron-based renderers.
The GS surpasses NeRF in terms of both training and inference speed, distinguishing itself by not depending on a neural network for operation. Rather, GS stores essential information within its 3D Gaussian components, a feature that renders it well-suited for dynamic scene modeling (Wu et al., 2023). Additionally, integrating GS with a dedicated 3D computer graphics engine is a straightforward process (Kerbl et al., 2023). However, conditioning GS is a challenging task due to the large number of Gaussian components typically involved, which can reach hundreds of thousands.
GaussianAvatars (Qian et al., 2023), utilizes the local coordinate system to generate Gaussians corresponding to the mesh’s faces. This approach is specifically designed for avatars and assumes the availability of a realistic (external) model for mesh fitting. However, training both the mesh and GS simultaneously is not feasible. While these solutions offer certain advantages, they do not directly combine Gaussian components with the mesh. As a result, automatic adaptation of the Gaussian parameters to changing mesh is impossible.
Figure 3: Visualization of affine transformation of Gaussian components when we modify mesh. Mesh vertices parameterize the mean and covariance matrix of Gaussian. Therefore, such parameters update when we change the mesh.
The more advanced method is proposed by SuGaR (Guédon & Lepetit, 2023), which introduces a regularization term in the GS cost function to encourage alignment between the Gaussians and the scene surface. SuGaR achieves this by using signed distance functions (SDF) and minimizing the difference between the SDF and the computed Gaussian values. Furthermore, the authors propose a method for extracting mesh from the GS model and introduce an optional refinement strategy that binds Gaussians to the mesh surface and jointly optimizes these Gaussians and the mesh through GS rendering.
Our work shows we can produce an editable GS model without expensive preprocessing stages like those used in SuGaR.
3 GaMeS : Mesh-Based Gaussian Splatting
Here, we delve into the details of the GaMeS model, commencing with the fundamental aspects of vanilla GS. Subsequently, we elucidate how we parameterize Gaussian distributions on the mesh faces. Then, we investigate extracting a mesh face from a single Gaussian component. Such face is used to reparametrize Gaussian components to be editable. Finally, we introduce the novel GaMeS approach.
Gaussian Splatting
The GS technique captures a 3D scene through an ensemble of 3D Gaussians, each defined by its position (mean), covariance matrix, opacity, and color. Additionally, spherical harmonics (SH) are employed to depict the color attributes. (Fridovich-Keil et al., 2022; Müller et al., 2022). The GS effectiveness is primarily attributed to the rendering process, which utilizes projections of Gaussian components. GS relies on a dense set of 3D Gaussian components with color and opacity:
$\mathcal{G}=\{(\mathcal{N}({\bf m}_{i},\Sigma_{i}),\sigma_{i},c_{i})\}_{i=1}^{n},$
where ${\bf m}_{i}$ denotes position, $\Sigma_{i}$ marks covariance, $\sigma_{i}$ stands for opacity, and $c$ is SH colors of $i$-th component.
In practice, GS uses factorization $\Sigma=R^{T}SSR$. We can train GS with a flat Gaussian distribution by forcing $s_{1}=0$, where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$.
The parameters of Gaussian distributions undergo direct training through gradient optimization. To enhance adaptability to complex scenes, GS employs additional training strategies. Notably, significant Gaussian components are subdivided. If the update parameters are substantial, Gaussians components are duplicated, which also can be removed due to their low transparency. In addition, the GS training procedure is implemented in the CUDA kernel, which supports fast training and real-time rendering.
Figure 4: The left image presents the Gaussian component constructed on the face $V=\{{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}\}$. We overlap the centers of Gaussian and face and parameterize the covariance matrix by vertices $V$. The right image presents how our model uses previously estimated Gaussian to construct $k$ (in Fig. $k=3$) components in GS. GaMeS components are built by scaling and a translation on a mesh face.
Distribution on mesh faces
In GaMeS, we placed all Gaussian components onto the mesh surface. Let us examine a single triangle face with vertices:
$V=\{{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}\}\subset\mathbb{R}^{3}.$
We aim to parameterize the Gaussian components using vertices from the face $V$. We express the mean vector as a convex combination of vertices $V$, thus determining the Gaussian splats positions:
${\bf m}_{V}(\alpha_{1},\alpha_{2},\alpha_{3})=\alpha_{1}{\bf v}_{1}+\alpha_{2}{\bf v}_{2}+\alpha_{3}{\bf v}_{3},$
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are trainable parameters such that $\alpha_{1}+\alpha_{2}+\alpha_{3}=1$. Through this parameterization, we consistently maintain the Gaussians positioned in the middle of the face $V$.
Figure 5: GaMeS also works when we do not have a mesh. In such a situation, we estimate pseudo-mesh obtaining a set of disconnected faces. Our pseudo-mesh is dedicated to modifying the GS model, not approximating objects’ surfaces.
Original Gaussian Create face Parameterized Gaussian
Figure 6: Each Gaussian corresponds to one face from the pseudo-mesh. The model parameterizes the Gaussian depending on the vertices, allowing it to be edited.
Fig. 3 shows examples of such transformation for rotation and scaling. Face $V$ at time zero corresponds to the triangle $V=\left\{{\bf v}^{V}_{1},{\bf v}^{V}_{2},{\bf v}^{V}_{3}\right\}$, then it is transformed into $V^{\prime}=\left\{{\bf v}^{\prime V}_{1},{\bf v}^{\prime V}_{2},{\bf v}^{\prime V}_{3}\right\}$. Three Gaussians splats $\mathcal{G}_{V}^{1}$, $\mathcal{G}_{V}^{2}$, $\mathcal{G}_{V}^{3}$ depend only on the position of the vertices.
The covariance matrix can be defined as an empirical covariance calculated from three points. However, such a solution is complex to combine with the optimization proposed in GS. Instead, the covariance is parameterized by factorization:
$\Sigma=R^{T}SSR,$
where $R$ is the rotation matrix and $S$ the scaling parameters. Here, we define rotation and scaling matrices to stay in the original framework. Let start from orthonormal vectors: ${\bf r}_{1},{\bf r}_{2},{\bf r}_{3}\in\mathbb{R}^{3}$, consisted into rotation matrix $R_{V}=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}].$ Where the first vector is defined by the normal vector:
${\bf n}=\frac{({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})\|},$
where $\times$ is the cross product. Given an explicitly defined mesh, we consistently possess knowledge of the vertex order for any given face. Hence, to calculate ${\bf r}_{2}$, we define the vector from the center to the vertex ${\bf v}_{1}$:
${\bf r}_{2}=\frac{{\bf v}_{1}-{\bf m}}{\|{\bf v}_{1}-{\bf m}\|},$
where ${\bf m}=\mathrm{mean}({\bf v}_{1},{\bf v}_{2},{\bf v}_{3})$, which corresponds to the centroid of the triangle.
The last operation is obtained through orthonormalizing the vector with respect to the existing two vectors (a single step in the Gram–Schmidt process (Björck, 1994)):
$\mathrm{orth}({\bf x};{\bf r}_{1},{\bf r}_{2})=x-\mathrm{proj}({\bf x},{\bf r}_{1})-\mathrm{proj}({\bf x},{\bf r}_{2}),$
where
$\mathrm{proj}({\bf v},{\bf u})=\frac{\langle{\bf v},{\bf u}\rangle}{\langle{\bf u},{\bf u}\rangle}{\bf u}.$
To obtain ${\bf r}_{3}$, we use the vector from the center to the second vertex of the triangle:
${\bf r}_{3}=\frac{\mathrm{orth}({\bf v}_{2}-{\bf m};{\bf r}_{1},{\bf r}_{2})}{\|\mathrm{orth}({\bf v}_{2}-{\bf m};{\bf r}_{1},{\bf r}_{2})\|}.$
In consequence, we obtain a rotation matrix $R_{V}=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, which aligns with the triangular face. As scaling parameter $S$ we use:
$S_{V}=\mathrm{diag}(s_{1},s_{2},s_{3}),$
where $s_{1}=\varepsilon,$ $s_{2}=\|\mathrm{mean}({\bf v}_{1},{\bf v}_{2},{\bf v}_{3})-{\bf v}_{1}\|,$ and $s_{3}=\langle{\bf v}_{2},{\bf r}_{3}\rangle.$ The first scaling parameters correspond with the normal vector. Since, in our case, the Gaussians are positioned on the faces, aligning with the surface $s_{1}$ should be equal to zero. To avoid numerical problems, we fix this value as a small value (constant).
Original mesh Subdivide larger faces
Refer to caption Refer to caption
Figure 7: GaMeS use a fixed number of faces. In the case of a mesh containing faces of different sizes, finding such parameters causes problems. In such a case, we propose to divide large faces into smaller parts. The left side includes only the original mesh, while on the right, the sizable mesh faces of the brown pool surrounding have been subdivided into smaller ones.
Only Gaussians means are on mesh   GaMeS      Only Gaussians means are on mesh     GaMeS
Refer to caption Refer to caption
Figure 8: GaMeS parameterizes all Gaussian components by mesh vertices to lie directly on the mesh. Therefore, when we modify the mesh positions, we automatically adapt the components’ means and covariances. Consequently, we can see differences in the rendered image quality for the transforming when (a) only the Gaussian position is editable and (b) when the Gaussian rotation and scaling are editable.
Consequently, the covariance of Gaussian distribution positioned to face is given by:
$\Sigma_{V}=R_{V}^{T}S_{V}S_{V}R_{V},$
and correspond with the shape of a triangle $V$. Fig.  4 shows the process of determining Gaussians. Here the $\alpha_{j}^{i}$ refer to the $i$-th Gaussian $\mathcal{G}_{V}^{i}$ and the $j$-th vertex, respectively.
Within our model, we train a scale parameter, denoted as $\rho$, to dynamically adjust the Gaussian component size. For one face let us take $k\in\mathbb{N}$ number of Gaussians:
$\mathcal{G}_{V}=\{\mathcal{N}({\bf m}_{V}(\alpha^{i}_{1},\alpha^{i}_{2},\alpha^{i}_{3}),\rho^{i}\Sigma_{V}))\}_{i=1}^{k},$
where $\alpha^{i}_{1}+\alpha^{i}_{2}+\alpha^{i}_{3}=1$ and $\rho^{i}\in\mathbb{R}_{+}$.
This distribution method enables us to render Gaussian splats solely dependent on the position of the mesh triangles. Consequently, we can readily perform geometric transformations when dealing with Gaussians within a single (triangle) face. Triangle transformation will apply Gaussian transformation as well.
Distribution on pseudo-mesh faces
When we do not have mesh or any initialization, we can automatically produce pseudo-mesh. First, we train GS with flat Gaussian’s using $s_{1}=\epsilon$ in training. Then, we estimate one triangle face for each Gaussian component and parameterize Gaussian components by the vertices of the mesh face, see Fig. 5. After reparametrization, we have final GaMeS.
We use one pseudo-mesh face for each Gaussian component. Let us assume that we have a Gaussian component parameterized by mean ${\bf m}$, rotation matrix $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$ and scaling $S=\mathrm{dig}(\varepsilon,s_{2},s_{3}).$ We define three vertex of our pseudo-mesh face, see, Fig. 6
$V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}],$
where ${\bf v}_{1}={\bf m}$, ${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$.
Now we froze the vertex of the face $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$ and reparameterize the Gaussian component by defining $\hat{\bf m}$, $\hat{R}=[\hat{\bf r}_{1},\hat{\bf r}_{2},\hat{\bf r}_{3}]$ and $\hat{S}=\mathrm{dig}(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}).$ First, we put $\hat{\bf m}={\bf v}_{1}$. The first vertex of $\hat{R}$ is given by a normal vector:
$\hat{\bf r}_{1}=\frac{({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})\|},$
where $\times$ is the cross product. The second one is defined by $\hat{\bf r}_{2}=\frac{({\bf v}_{2}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\|}.$ The third one is obtained as a single step in the Gram–Schmidt process (Björck, 1994):
$\hat{\bf r}_{3}=\mathrm{orth}({\bf v}_{3}-{\bf v}_{1};{\bf r}_{1},{\bf r}_{2}).$
Scaling parameters can also be easily calculated as $s_{1}=\varepsilon$, $\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|$ and $\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle$.
Consequently, the covariance of Gaussian distribution positioned to face is given by:
$\hat{\Sigma}_{V}=\hat{R}_{V}^{T}\hat{S}_{V}\hat{S}_{V}\hat{R}_{V},$
and correspond with the shape of a triangle $V$. For one face $V$, we define the corresponding Gaussian component:
$\hat{\mathcal{G}}_{V}=\mathcal{N}(\hat{\bf m}_{V},\hat{\Sigma}_{V})).$
We obtain the Gaussian component parameterized by face mesh. We can use such a strategy GS with a flat Gaussian component.
GaMeS: Gaussian Splatting on Mesh
The essential step in our model involves positioning Gaussians on the 3D mesh of the scene or object in question. Here, we have to consider two crucial experimental scenarios, namely, with and without using mesh during the training. First, we assume that a mesh is being provided and exclusively train the Gaussian components’ centers, colors, and opacity. Here, numerous methods facilitate the generation of an initial mesh for objects, but our primary focus is not on this task. Instead, we emphasize addressing distinct issues that extend beyond the scope of initializing object meshes like EG3D (Chan et al., 2022), NGLOD (Takikawa et al., 2021b) or NeRFMeshing (Rakotosaona et al., 2023).
Second, we assume that we do not have a mesh. In the latter case, initialization of the mesh is required and achieved by utilizing our strategy with extracting pseudo-mesh from pretrained GS.
Additionally, we also show the initialization of the mesh using Faces Learned with an Articulated Model and Expressions (FLAME) (Li et al., 2017). The mesh initialization of the human face is not perfected fitted. We adjust the mesh vertices and Gaussian parameters throughout the training process.
Let us consider a mesh denoted as $\mathcal{M}=(\mathcal{V},\mathcal{E}),$ where $\mathcal{V}\subset\mathbb{R}^{3}$ represents vertices, and $\mathcal{E}\subset\mathbb{R}^{3}$ signifies edges. The mesh’s faces, denoted by =$\mathcal{F}_{\mathcal{M}}=\left\{\left({\bf v}^{V_{i}}_{1},{\bf v}^{V_{i}}_{2},{\bf v}^{V_{i}}_{3}\right)\right\}_{i}^{n}$ are described by sets of vertices. As previously detailed, such mesh serves as a parameterization for shaping Gaussian components.
For our experiments, we choose the number $k\in\mathbb{N}$ of Gaussian components per mesh face and define how densely they should be placed on its surface. For mesh containing $n\in\mathbb{N}$ faces, the final number of Gaussians is fixed and equal to $k\cdot n$.
In the scenario with known mesh, for each Gaussian component, we define trainable parameters used to calculate mean and covariance $\{(\alpha^{i}_{1},\alpha^{i}_{2},\alpha^{i}_{3},\rho^{i})\}_{i=1}^{k\cdot n}$. Therefore, GaMeS use a dense set $\mathcal{G}_{\mathcal{M}}$ of 3D Gaussians:
$\mathcal{G}_{\mathcal{M}}\!\!=\!\!\!\bigcup_{V_{i}\in\mathcal{F}_{\mathcal{M}}}\{\mathcal{G}_{V_{i}},\sigma_{i,j},c_{i,j}\}_{j=1}^{k},$
where $\sigma_{i,j}$ is opacity, and $c_{i,j}$ is SH colors of $j$-th Gaussian, and $i$-th face.
When we do not have mesh, we train GS, and for each Gaussian, we estimate face. Then we obtain set $\mathcal{G}_{\mathcal{M}}$ of 3D parameterized Gaussians $\hat{\mathcal{G}}_{V}$ on pseudo-mesh $\hat{\mathcal{M}}$:
$\mathcal{G}_{\mathcal{M}}\!\!=\!\!\!\bigcup_{V_{i}\in\mathcal{F}_{\hat{\mathcal{M}}}}\{\hat{\mathcal{G}}_{V_{i}},\sigma_{i},c_{i}\},$
where $\sigma_{i}$ is opacity, and $c_{i}$ is SH colors Gaussian corresponded to $i$-th face.
Mesh initialization
During training, we can modify mesh vertices and Gaussian parameters. We can also use external tools for mesh fitting (Takikawa et al., 2021a) and train only Gaussian components. In the experimental section, we show both of such scenarios.
In our model, we use a fixed number of Gaussian components per face. In the case of models with various face sizes, finding such parameters is not trivial. To solve such a problem, we propose to divide large faces into smaller parts, see Fig. 7.
PSNR $\uparrow$
Chair Drums Lego Mic Materials Ship Hotdog Ficus
Static
NeRF 33.00 25.01 32.54 32.91 29.62 28.65 36.18 30.13
VolSDF 30.57 20.43 29.46 30.53 29.13 25.51 35.11 22.91
Ref-NeRF 33.98 25.43 35.10 33.65 27.10 29.24 37.04 28.74
ENVIDR 31.22 22.99 29.55 32.17 29.52 21.57 31.44 26.60
Plenoxels 33.98 25.35 34.10 33.26 29.14 29.62 36.43 31.83
Gaussian Splatting 35.82 26.17 35.69 35.34 30.00 30.87 37.67 34.83
Editable
GaMeS (our) 35.38 26.03 35.89 37.16 29.62 31.55 37.56 35.12
GaMeS (our - pseudo-mesh) 34.74 25.94 35.88 36.58 30.50 30.66 37.84 35.41
RIP-NeRF 34.84 24.89 33.41 34.19 28.31 30.65 35.96 32.23
Table 1: Quantitative comparisons (PSNR) on a NeRF-Synthetic dataset showing that GaMeS gives comparable results with other models.
Initial Position GaMeS Only Gaussians means on mesh Refer to caption
Figure 9: Compared to a method solely based on Gaussian averaging dependent on position, we observe that, following rotation and stretching the surface, Gaussians should dynamically adjust their scale and rotation to seamlessly conform to the change as in the case of GaMeS model.
Mesh modification
GaMeS parameterize all Gaussian components by verities of mesh in such a way that they lie on meshes. Therefore, when we modify mesh positions, we automatically adapt the means and covariances of relevant Gaussians. Such an approach allows us to work with transformation. To better visualize such properties, we compare our model with a solution where the Gaussian components’ centers are placed on mesh faces but do not parameterize the covariance matrix. In Fig. 8 we can see that GaMeS Gaussian component adapts automatically to mesh modification.
In particular, when an object is bent, models that would only consider the position (means) of Gaussians depending on the mesh would fail to adapt properly. We can observe this effect in Fig. 8 with the ship’s rope rendering. In contrast, GaMeS provides a perfect fit with the modification, and the operation scheme is shown in Fig. 9.
4 Experiments
Here, we delineate implementation details, describe the used datasets, and the reasons for their selection. Then, we describe three scenarios: (1) when we have an existing mesh, (2) when we use pseudo-mesh estimated during training, and (3) when we use FLAME initialization. In the end, we show inference time both on static and animated scenes.
4.1 Implementation details and data description
                              PSNR $\uparrow$
Outdoor scenes Indoor scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Static
Plenoxels 21.91 20.10 23.49 20.66 22.25 27.59 23.62 23.42 24.66
INGP-Big 22.17 20.65 25.07 23.47 22.37 29.69 26.69 29.48 30.69
Mip-NeRF360 24.37 21.73 26.98 26.40 22.87 31.63 29.55 32.23 33.46
GS - 7K 23.60 20.52 26.25 25.71 22.09 28.14 26.71 28.55 28.85
GS - 30K 25.25 21.52 27.41 26.55 22.49 30.63 28.70 30.32 31.98
Editable
R-SuGaR-15K 23.14 - 25.36 24.70 - 30.03 27.62 29.56 30.51
GaMeS (Our-pseudo-mesh) 24.99 21.27 27.22 26.54 22.39 31.52 28.92 31.12 32.09
Table 2: The quantitative comparisons of reconstruction capability (PSNR) on Mip-NeRF360 dataset. R-SuGaR-15K uses 1M vertices. GaMeS uses pseudo-mesh.
The time taken by the optimization process for our model varied from a few up to no more than sixty minutes, as it depends on factors like scene complexity, dataset characteristics, the number of Gaussians per face, and vertex fine-tuning decisions. Nonetheless, all experiments were conducted within a reasonable time frame, underscoring the efficiency of our approach across diverse scenarios. The supplementary material provides all the relevant details, and the source code is available in Github [^1] https://github.com/waczjoan/gaussian-mesh-splatting.
We illustrated the fundamental gains of GaMeS model through experiments with three distinct datasets.
Synthetic Dataset: A NeRF-provided eight geometrically complex objects with realistic non-Lambertian materials (Mildenhall et al., 2020). In addition to containing images of the objects, the dataset also incorporates corresponding meshes. We leveraged this dataset for our initial experiments to unequivocally showcase the capabilities of our model and underscore its flexibility in mesh editing.
Mip-NeRF360 Dataset: A dataset comprising five outdoor and four indoor scenes, each featuring intricate central objects or areas against detailed backgrounds (Barron et al., 2022). We used this dataset to demonstrate the adaptability of our model in handling dense scenes, employing a non-conventional strategy during the initial training stage in mesh preparation.
Human Faces Dataset: A modified subset of the Stirling/ESRC 3D Face Database [^2] https://pics.stir.ac.uk/ESRC/index.htm, that includes images of six people, generated using Blender software from various perspectives, consistently excluding the backs of heads. Only one expression is assigned for each face. Notably, the dataset lacks corresponding mesh representations for the depicted person. This dataset is used to compare GaMeS with the NeRFlame model (Zając et al., 2023), mainly within the human-mesh fitting task.
4.2 Scenario I: Model with existing mesh
In this scenario, we utilize the provided mesh, incorporating Gaussians by strategically placing them onto its surface. We also allowed for moving vertices. Our initial experiments used the Synthetic Dataset (Mildenhall et al., 2020), incorporating shared meshes. Table 1 presents PSNR metrics demonstrating the competitiveness of GaMeS approach with existing methods. Additional numerical comparisons are available in the supplementary material. Notably, a significant contribution lies in the ease with which renders can be modified manually or in an automated process. Fig. 1 shows examples of reconstruction and simple object animations.
Figure 10: The main idea of fitting human mesh and how it can be transformed into a different person.
4.3 Scenario II: A model without mesh
When we do not have mesh, we train GS with flat Gaussians. Then, we construct the pseudo-mesh. Table 1 shows a comparison with other models on the Synthetic Dataset. Our approach is also capable of modeling unbounded scenes. In order to achieve that, we first create an initial pseudo-mesh and then adjust it during the training. Results in Table 2 demonstrate that such modeling is feasible and reasonably comparable to other existing models. Particularly, the results show parity with the GS-30K and R-SuGaR-15K models.
As anticipated, we observe scene reconstruction, including both objects and details, as can be seen in Fig. 2. Nevertheless, backgrounds and elements with large mesh faces may exhibit a lower-quality appearance. However, this does not hinder the feasibility of GaMeS model for reconstruction and modification.
4.4 Scenario III: GaMeS with FLAME as Init Mesh
Figure 11: Reconstruction of 3D peoples and generation of new expressions using the GaMeS model.
In this experiment, we show the possibility of fitting mesh using an Init Mesh, which was not acquired directly from our data in contrast to previous scenarios. Here, we relied on the Faces Dataset as it provides such required experimental data. To initialize the meshes, we used the FLAME (Li et al., 2017) framework, generating fully controllable meshes of human faces. Consequently, we could highlight the key advantage of GaMeS model, i.e., the ability to modify objects. We used the official implementation of the FLAME, with the number of parameters suggested in RingNet (Sanyal et al., 2019) In addition, we implemented an extra detail learning parameter to better represent hair or T-shirts.
In Fig. 10, we show the main idea behind our approach. Initially, a human mesh is generated using the FLAME model. Thanks to accurately adjusting the mesh, we can apply it, for example, to another person. By utilizing the FLAME parameters, we can modify various aspects, especially facial expressions, to generate facial expressions like smiling, sadness, eye closure, mouth opening, and flirty or disgusted expressions. The possibilities of reconstruction and modifications for all six faces are depicted in Fig. 11. All expressions were chosen randomly as our main goal was to show GaMeS model’s ability to render objects as well as its modifications/animations.
PSNR $\uparrow$
Person 1 2 3 4 5 6
Static
NeRF 33.37 33.39 33.08 31.96 33.15 32.42
GS 50.16 49.04 49.81 49.97 49.77 49.17
Editable
NeRFlame 27.89 29.79 29.70 25.78 32.59 29.18
GaMeS 32.73 29.56 29.42 29.95 32.27 31.50
Table 3: Comparison of PSNR, SSIM, and LPIPS matrices between static models: NeRF, GS and referring to them: NeRFlame and GaMeS. We used 100 Gaussians per face.
The comparison of the mesh fitting capability between the GaMeS with the NeRFlame model is shown in Fig. 12. Notably, the faces are distinctly delineated and closely resemble the modeled objects, capturing not only postures but also expressions and intricate details. Table 3 shows the results of using PSNR metrics on six faces. Comparisons are made between the two baseline NeRF models, GS, and the models created from them, i.e., the corresponding NeRFlame and GaMeS with animation capabilities.
Figure 12: During the training of GaMeS, we simultaneously model FLAME mesh. Here, we present the meshes fitted by GaMeS compared to fitted mesh by NeRFlame model.
5 Conclusion
This paper presents a new GS-based model called GaMeS that leverages two main concepts, i.e., classical mesh and vanilla GS. Here, we represent a 3D scene with a set of Gaussian components lying on the mesh. We can use the initial mesh as an input or estimate it during training. Moreover, we parameterize the Gaussian components by the mesh vertices, allowing single-stage training. In inference, we obtain GS, which can be modified in real-time by manipulating mesh components. As a result, when the mesh changes, all Gaussian components automatically adapt to a new shape.
Limitations GaMeS allows real-time modifications, but artifacts appear in the case of significant changes in the case of meshes with large faces. In practice, large faces should be divided into smaller ones. How to change Gaussian components in GaMeS when mesch faces are split is not apparent.
Societal impact Our GaMeS contributes to lowering the computational complexity of 3D modeling, therefore reducing the energy consumption of such models.
Appendix A Appendix
The appendix to the paper contains excerpts related to NeRF Synthetic and a continuation of the numerical results expounded in the main document.
PSNR $\uparrow$
Number of Gaussians Chair Drums Lego Mic Materials Ship Hotdog Ficus
1 29.32 25.18 32.74 32.48 25.37 27.45 32.10 Z32.05
5 30.70 25.35 33.21 32.49 25.29 29.29 33.40 32.29
10 31.11 25.36 OOM 32.29 25.25 29.56 33.65 32.35
pseudo-mesh 35.34 25.93 35.42 35.34 30.01 29.58 37.65 34.81
SSIM $\uparrow$
Chair Drums Lego Mic Materials Ship Hotdog Ficus
1 0.934 0.940 0.972 0.984 0.916 0.879 0.959 0.978
5 0.950 0.942 0.975 0.984 0.914 0.894 0.968 0.978
10 0.951 0.942 OOM 0.983 0.913 0.896 0.970 0.979
pseudo-mesh 0.986 0.952 0.982 0.991 0.958 0.886 0.985 0.987
LPIPS $\downarrow$
Chair Drums Lego Mic Materials Ship Hotdog Ficus
1 0.066 0.052 0.027 0.012 0.060 0.133 0.062 0.018
5 0.050 0.049 0.022 0.012 0.061 0.107 0.042 0.018
10 0.044 0.049 OOM 0.012 0.063 0.102 0.038 0.019
pseudo-mesh 0.012 0.039 0.015 0.006 0.033 0.105 0.019 0.012
Table 4: The quantitative comparisons (PSNR / SSIM / LPIPS) on Synthetic dataset with original meshes or pseudo-mesh, with white background. OOM - CUDA out of memory. In this experiment we used Tesla V100-SXM2-32GB GPU.
A.1 Number of Gaussians per face
Here, we show the results of the experiments for the NeRF Synthetic dataset with a white background and original mesh. The dependence of the PSNR, SSIM and LPIPS result on the number of Gaussians per face is shown in Table 4.
Mostly, as a consequence of increasing the number of Gaussians, we get better results, and we can observe that a single Gaussian per face is insufficient. Materials and Lego has a very dense mesh, consequently positioning an excessive number of Gaussians within a confined space does not yield substantial improvements in the results.
Initial Mesh Subdivided Mesh
$k=1$ $k=10$ $k=5$
PSNR = 29.32 PSNR = 31.11 PSNR = 33.01
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: A qualitative assessment delineating the impact of varying Gaussian quantities per face, discerned between the initial mesh and its subdivided counterpart.
Initial Mesh Subdivided Mesh k=1 k=10 k=5
Figure 14: A juxtaposition of outcomes is conducted for a model varying in Gaussian counts per face and with a mesh subdivided. GaMeS with good mesh initialization is able to capture fine details.
However, experiments have shown that choosing a fixed number of Gaussians per face mesh is inefficient when the mesh contains different face sizes – in particular, it is difficult to cover large faces. Therefore, we decided to split the large facets. Splitting large facets, even using fewer Gaussians per face, allows for significantly better results and detail capturing. The effects are shown in Fig. 13 (initial positions) and Fig. 14 (modifications).
A.2 Modification
Illustrating the reconstruction and modification through the application of the GaMeS model is shown in Fig. 15. GaMeS allows the creation of lifelike alterations, such as elevating the excavator blade or imaginative and surrealistic bends.
In fact, modification depends only on user skills. Fig. 16 shows various modifications of the ficus such as: a) scaling b) modifying a certain part – like bending branches, or modifying the whole object – dancing ficus. (c) ignoring a certain part. All images are generated from the view of the same camera. Note that Gaussians represent color from two sides. Therefore, after removing the soil from the pot, we can still see the inner side of the pot.
Fig. 17 shows that modifications are possible even using only pseudo-mesh.
Before modification   After modification
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 15: An example of reconstruction and modification using the GaMeS model. The model enables more realistic modifications like lifting the excavator blade as well as surrealistic drums bending.
Figure 16: Demonstrations of potential modifications facilitated by games include: a) scaling, b) transforming either specific components or entire objects, and c) selectively disregarding specific sections.
Figure 17: An example of modification using the GaMeS model with pseudo-mesh.
A.3 Extension of numerical results from main paper
Here, we present extensions of experiments proposed in the main part.
PSNR $\uparrow$
Outdoor scenes Indoor scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Static
Plenoxels 21.91 20.10 23.49 20.66 22.25 27.59 23.62 23.42 24.66
INGP-Big 22.17 20.65 25.07 23.47 22.37 29.69 26.69 29.48 30.69
Mip-NeRF360 24.37 21.73 26.98 26.40 22.87 31.63 29.55 32.23 33.46
GS - 7K 23.60 20.52 26.25 25.71 22.09 28.14 26.71 28.55 28.85
GS - 30K 25.25 21.52 27.41 26.55 22.49 30.63 28.70 30.32 31.98
Editable
R-SuGaR-15K 22.91 - 25.29 24.55 - 29.95 27.47 29.38 30.42
GaMeS (Our-pseudo-mesh) 24.99 21.27 27.22 26.54 22.39 31.52 28.92 31.12 32.09
SSIM $\uparrow$
Outdoor scenes Indoor scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Static
Plenoxels 0.496 0.431 0.606 0.523 0.509 0.842 0.759 0.648 0.814
INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906
Mip-NeRF360 0.685 0.583 0.813 0.744 0.632 0.913 0.894 0.920 0.941
GS - 7K 0.675 0.525 0.836 0.728 0.598 0.884 0.873 0.900 0.910
GS - 30K 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938
Editable
R-SuGaR-15K 0.631 - 0.771 0.681 - 0.909 0.890 0.907 0.933
GaMeS (Our-pseudo-mesh) 0.757 0.596 0.860 0.766 0.628 0.916 0.904 0.923 0.939
LPIPS $\downarrow$
Outdoor scenes Indoor scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Static
Plenoxels 0.506 0.521 0.386 0.503 0.540 0.419 0.441 0.447 0.398
INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205
Mip-NeRF360 0.301 0.344 0.170 0.261 0.339 0.211 0.204 0.127 0.176
GS - 7K 0.318 0.417 0.153 0.287 0.404 0.272 0.254 0.161 0.244
GS - 30K 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205
Editable
R-SuGaR-15K 0.349 - 0.218 0.336 - 0.243 0.234 0.166 0.219
GaMeS (Our-pseudo-mesh) 0.216 0.342 0.111 0.219 0.332 0.221 0.203 0.129 0.205
Table 5: The quantitative comparisons (PSNR / SSIM / LPIPS) on the Mip-NeRF360 dataset. R-SuGaR-15K, with the number of 200K vertices.
PSNR $\uparrow$
Chair Drums Lego Mic Materials Ship Hotdog Ficus
Static
NeRF 33.00 25.01 32.54 32.91 29.62 28.65 36.18 30.13
VolSDF 30.57 20.43 29.46 30.53 29.13 25.51 35.11 22.91
Ref-NeRF 33.98 25.43 35.10 33.65 27.10 29.24 37.04 28.74
ENVIDR 31.22 22.99 29.55 32.17 29.52 21.57 31.44 26.60
Plenoxels 33.98 25.35 34.10 33.26 29.14 29.62 36.43 31.83
Gaussian Splatting 35.82 26.17 35.69 35.34 30.00 30.87 37.67 34.83
Editable
GaMeS (our) 35.38 26.03 35.89 37.16 29.62 31.55 37.56 35.12
GaMeS (our - pseudo-mesh) 34.74 25.94 35.88 36.58 30.50 30.66 37.84 35.41
RIP-NeRF 34.84 24.89 33.41 34.19 28.31 30.65 35.96 32.23
SSIM $\uparrow$
Static
NeRF 0.967 0.925 0.961 0.980 0.949 0.856 0.974 0.964
VolSDF 0.949 0.893 0.951 0.969 0.954 0.842 0.972 0.929
Ref-NeRF 0.974 0.929 0.975 0.983 0.921 0.864 0.979 0.954
ENVIDR 0.976 0.930 0.961 0.984 0.968 0.855 0.963 0.987
Plenoxels 0.977 0.933 0.975 0.985 0.949 0.890 0.980 0.976
Gaussian Splatting 0.987 0.954 0.983 0.991 0.960 0.907 0.985 0.987
Editable
GaMeS (our) 0.987 0.953 0.982 0.992 0.952 0.904 0.985 0.986
GaMeS (our - pseudo-mesh) 0.983 0.945 0.981 0.992 0.958 0.888 0.985 0.986
RIP-NeRF 0.980 0.929 0.977 0.962 0.943 0.916 0.963 0.979
LPIPS $\downarrow$
NeRF 0.046 0.091 0.050 0.028 0.063 0.206 0.121 0.044
VolSDF 0.056 0.119 0.054 0.191 0.048 0.191 0.043 0.068
Ref-NeRF 0.029 0.073 0.025 0.018 0.078 0.158 0.028 0.056
Plenoxels 0.031 0.067 0.028 0.015 0.057 0.134 0.037 0.026
ENVIDR 0.031 0.080 0.054 0.021 0.045 0.228 0.072 0.010
Gaussian Splatting 0.012 0.037 0.016 0.006 0.034 0.106 0.020 0.012
Editable
GaMeS (our) 0.009 0.038 0.014 0.005 0.042 0.090 0.017 0.012
GaMeS (our - pseudo-mesh) 0.012 0.039 0.016 0.006 0.036 0.109 0.018 0.012
RIP-NeRF - - - - - - - -
Table 6: Quantitative comparisons (PSNR) on a NeRF-Synthetic dataset showing that GaMeS gives comparable results with other models.